OpenStudy (anonymous):

an astronomical clock has 24 dials on it it the hour hand covers a complete revolution in 24 hour and the minute hand in one hour a) how many times are they perpendicular to each other in one hour b) how many times are they opposite to each other in 24 hour

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

@ganeshie8

OpenStudy (ikram002p):

maby in one hour 4 times ?

OpenStudy (ikram002p):

and b in 24 hours is 4*24 ??

OpenStudy (ikram002p):

im assuming that for a normal clock

OpenStudy (anonymous):

no

OpenStudy (anonymous):

http://www.orloj.com/orloj_dial_sphere.jpg

OpenStudy (dan815):

:)

OpenStudy (ikram002p):

i knw nothing abt this shape lol exept im leo :P

OpenStudy (anonymous):

does anyone know it

OpenStudy (anonymous):

virgo

OpenStudy (dan815):

write vector functions of hour hand wrt to time and samething for min hand

OpenStudy (anonymous):

dont know any of that

OpenStudy (dan815):

and solve for all time values where h'(t)=-1/min'(t)

OpenStudy (dan815):

kk well then lets use simple math

OpenStudy (dan815):

in 1 hour it will be perpendicular twice right

OpenStudy (dan815):

so in 24 hrs u will see it perpendicular 48 times

OpenStudy (ikram002p):

thats it ?

OpenStudy (dan815):

wait it only asks you perpendicular for 1 hour though

OpenStudy (dan815):

so ya thats it

OpenStudy (anonymous):

are you sure

OpenStudy (dan815):

now how many times would it be opposite each other in 24 hrs

OpenStudy (dan815):

lol ganeshi

OpenStudy (anonymous):

ganeshi what is it your freaking me breo

OpenStudy (dan815):

what is this meaning of this wolfram!

ganeshie8 (ganeshie8):

wolfram is giving up.. so we need to count manually how many solutions will be there i gueess..

OpenStudy (ikram002p):

×_×

ganeshie8 (ganeshie8):

those equations look okay @dan815 @ikram002p ?

OpenStudy (ikram002p):

dnt ask me i knw nothing, i dnt knw this clock sys :o

OpenStudy (dan815):

complete failure!

ganeshie8 (ganeshie8):

Hours dial : 1 revolution every 24 hours, so 1/24 revolutiosn every hour : y1 = sin(pi/12t) Minutes dial : 1 revolution every 1hours, so 1 revolutiosn every hour : y1 = sin(2pit)

ganeshie8 (ganeshie8):

oh u want a vector function ?

OpenStudy (dan815):

that was the other way

OpenStudy (ikram002p):

why exams r tough these days -.-

OpenStudy (dan815):

i thought of the most trivial answers when he shut down my vector ideas :)

OpenStudy (dan815):

complete mistake

ganeshie8 (ganeshie8):

Hours : y1 = sin(pi/12t) i + cos(pi/12t) j Minutes : y2 = sin(2pit) i + cos(2pit) j

ganeshie8 (ganeshie8):

there are ur vector functions, if they're right, i guess we can simply solve for y1'*y2' = -1 ?

ganeshie8 (ganeshie8):

unless our reasoning is flawed... lol

OpenStudy (dan815):

it must be fine solving that for t=0 to 24

ganeshie8 (ganeshie8):

for part a) 0<=t<=1 right ?

ganeshie8 (ganeshie8):

or any other interval of length 1

OpenStudy (dan815):

oh yes

OpenStudy (dan815):

wait does t=0 to t=1 complete a full day in these functions of yours

ganeshie8 (ganeshie8):

t is in hours

ganeshie8 (ganeshie8):

revolutions per hour it is..

OpenStudy (dan815):

starting at 6:6 position on a 24 hour clock hour hand r(t)=<cos(2pit),sin(2pi t)> min hand r2(t)=<cos(24*24*2pi t),sin(24*24*2pi t)>

OpenStudy (dan815):

@ganeshie8 u think that works? 24* 24

OpenStudy (dan815):

no 24*

OpenStudy (dan815):

just 1 24

ganeshie8 (ganeshie8):

hour hand covers a complete revolution in 24 hours

ganeshie8 (ganeshie8):

hour hand r(t)=<cos(2pit),sin(2pi t)>

ganeshie8 (ganeshie8):

how does this complete 1 revolution in 24 hours ?

OpenStudy (dan815):

in t=0 to t=1

ganeshie8 (ganeshie8):

you also using unit as hour right ?

ganeshie8 (ganeshie8):

oh you're using unit as day ?

ganeshie8 (ganeshie8):

then it makes sense to meh.. .

OpenStudy (dan815):

r(t)=<cos(2pi t),sin(2pi t)> r2(t)=<cos(24*2pi t),sin(24*2pi t)> t=0 to t=1 ya for 1 fll day

ganeshie8 (ganeshie8):

yah this looks good.. .

ganeshie8 (ganeshie8):

but, wat does it mean to multiply two vectors ?

OpenStudy (dan815):

we wont multiply

ganeshie8 (ganeshie8):

r'(t) * r2'(t)

ganeshie8 (ganeshie8):

uhmm

OpenStudy (dan815):

we will get the slope functions both both and see all the time values they are perpendicular

ganeshie8 (ganeshie8):

r(t)=<cos(2pi t),sin(2pi t)> r2(t)=<cos(24*2pi t),sin(24*2pi t)> r'(t) = 2pi<-sin(2pit) , cos(2pit) > r2'(t)=48pi<-sin(24*2pi t),cos(24*2pi t)>

OpenStudy (dan815):

okay so for t from 0 to 1 r'(t) dot r2(t)==0 return t

OpenStudy (dan815):

lol

OpenStudy (dan815):

if only the computer was smart enough to understand that

OpenStudy (dan815):

man ive been learning about circuits and gates.. its so amazing i feel like i can build those little computers people are making on minecraft soon

ganeshie8 (ganeshie8):

oh we will take the dot product and set it equal to 0

OpenStudy (dan815):

yah

ganeshie8 (ganeshie8):

oh nice, digital electronics or vlsi ?

OpenStudy (dan815):

i dunno its just a intro course EECS

OpenStudy (dan815):

pretty much a copy of that course

ganeshie8 (ganeshie8):

ohh u must be having fun :) digital part is easy, but analog design is a bit tough :o

OpenStudy (dan815):

yeah

OpenStudy (dan815):

its fun

OpenStudy (dan815):

that prof on mit guy is really cool too

OpenStudy (dan815):

lets oslve this i want to see what it comes out to

ganeshie8 (ganeshie8):

Anant Agrawal... indian guy lol

OpenStudy (dan815):

yeah haha

OpenStudy (dan815):

best teacher!